OpenStudy (anonymous):

Solve the absolute value equation. 2|3x – 5| – 8 = 8

OpenStudy (anonymous):

2|3x – 5| – 8 = 8 2|3x – 5| = 0 |3x – 5| = 0 x = 5/3